QUESTION 9

  1. Consider the following binary search tree:

    50 15 62 20 58) 91 37 60 24

    If the root node, 50, is deleted, which node will become the new root?  

    A

    15

    B

    24

    C

    37

    D

    62

  

QUESTION 10

  1. In the following trees EXCEPT______, the left and right subtrees of any node have heights that differ by at most 1.

    A

    complete trees

    B

    perfect full trees

    C

    balanced binary trees

    D

    binary search trees

  

QUESTION 11

  1. A perfect full binary tree whose height is 5 has ______ nodes. (Note: the height of a tree is the number of the levels in the tree.)

    A

    5

    B

    15

    C

    31

    D

    32

  

QUESTION 12

  1. A complete binary tree with 20 nodes has a height of ________. (Note: the height of a tree is the number of the levels in the tree.)

    A

    3

    B

    4

    C

    5

    D

    6
Answer #1

(9) If the root node, 50, is deleted, which node will become the new root:

Answer: (c) 37

Description:

Find the immediate successor or predecessor in the tree and replace that node with deleting node. Here, immediate predecessor that is maximum of left subtree is 37.

(10) In the following trees EXCEPT______, the left and right subtrees of any node have heights that differ by at most 1:

Answer: (d) binary search trees

Description:

In binary search tree, the left and right subtrees of any node have heights that differ by more than 1 value.

(11) A perfect full binary tree whose height is 5 has ______ nodes. (Note: the height of a tree is the number of the levels in the tree.):

Answer: (c) 31

Description:

A Perfect Binary Tree whose height is h has total 2h - 1 node. For h = 5, nodes are 2h - 1 = 25 - 1 = 32 - 1= 31.

(12) A complete binary tree with 20 nodes has a height of ____(Note: the height of a tree is the number of the levels in the tree.):

Answer: (b) 4

Description:

A complete binary is a binary tree whose all levels are completely filled up except the last level. So, From level 0, 1, 2, 3 and up to 4 has 20 nodes.